OpenStudy (anonymous):

The ratio of smokers to nonsmokers on a particular flight was 2:3. Smoking passengers represented five more than 1/3 of all the passengers aboard. How many passengers were on that flight?

OpenStudy (anonymous):

The ratio of smokers to nonsmokers on a particular flight was 2:3 = P = S + (3S/2) Smoking passengers represented five more than 1/3 of all the passengers aboard. P = (S + 5)*3 where P = total passengers S = Smokers (3S/2) = Non-Smokers

OpenStudy (anonymous):

Solve the system of equations from there

OpenStudy (anonymous):

How did you get 3S/2 for non- smokers?

OpenStudy (anonymous):

if the ratio of smokers to non is 3:2 then there is 3 nons for every 2 smokers, or 3/2 nons for every smoker, or (3/2 * S)

OpenStudy (anonymous):

kk thank you